Little Johnny invested $10,000. He put part of it in a bank at 5 percent interest. He put the remainder in bonds which pay a 9 percent annual return. How much did he invest in each method (savings bank and bond) if his annual income from the two investments was $660?
Answer by stanbon(75887)

Little Johnny invested $10,000. He put part of it in a bank at 5 percent interest. He put the remainder in bonds which pay a 9 percent annual return. How much did he invest in each method (savings bank and bond) if his annual income from the two investments was $660?
-------------------------------------------------
Equation:
Quantity Eq::: x + y = 10000
Interest Eq::0.05x+0.09y = 660
------------------------
Multiply thru the 1st equation by 5:
Multiply thru the 2nd equation by 100:
5x + 5y = 50000
5x + 9y = 66000
--------------------
Subtract the 1st from the 2nd and solve for "y":
4y = 16000
y = $4000 (amount invested at 5%)
---
Since x+y = 10000, x = $6000 (amount invested at 9%)
==========================================================
Cheers,
Stan H.